OpenStudy (anonymous):

Find two consecutive whole numbers that square root of 27 lies between

OpenStudy (anonymous):

whts

OpenStudy (welshfella):

what are the 2 square numbers that come before and after 27?

OpenStudy (anonymous):

No clue lol I am terrible at math

OpenStudy (anonymous):

3

OpenStudy (welshfella):

square numbers are 1, 2^2 (4) , 3^3 (9) , 4^2 ( 16) and 5^2

OpenStudy (welshfella):

* that is 3^2

OpenStudy (welshfella):

what is 5^2?

OpenStudy (welshfella):

5^2 = 5*5 = ?

OpenStudy (anonymous):

25

OpenStudy (welshfella):

i'm talking to myself lol!

OpenStudy (welshfella):

oh some ones's listening yes

OpenStudy (anonymous):

Im so sorry lol I had to go do something for my brother keep explaining

OpenStudy (anonymous):

@welshfella

OpenStudy (anonymous):

@Hero @iGreen @perl do y'all think you can help me please I seriously don't get it

OpenStudy (zzr0ck3r):

5^2 = 25 6^2 = 36 \(5<\sqrt{27}<6\)

OpenStudy (anonymous):

omg thank you

OpenStudy (zzr0ck3r):

you should be able to do any problem similar to this now.